# Modeling Large Nonstationary Spatial Data with the Full-Scale Basis Graphical Lasso

### v1

@@ 3.4.5 Prediction
While the model was fit on data with a $2^{\text{o}}$ resolution in latitude and longitude, the model can be run at higher resolutions, for example we have limited access to $0.25^{\text{o}}$ resolution data. We consider this subgrid-scale data to test the predictive performance of the model by calculating the continuous rank probability score (CRPS) and the root mean square error (RMSE). Reported CRPS and RMSE values are averaged over individual predictions; CRPS has a closed form for Gaussian predictive distribution ([scoring]),
@@
From these results, it is clear that predictions are much more accurate using the FSBGL with the mean and median CRPS, as well as the RMSE, being less than half that of the BGL and LatticeKrig. In fact, the lowest CRPS obtained by LatticeKrig at any predicted location is 4.64, which is more than the median CRPS for any of the models fitted with the FSBGL. This suggests that the FSBGL is also suitable to be used to fit predictive distributions to data exhibiting multiscale behavior, for example one potential route of future investigation could be whether this technique can be allow for running climate models at lower spatial resolution from which FSBGL can statistically downscale to a fine resolution. This should be especially contrasted with the full-scale approximations of [sangfullscale], for example, as the reliance on EOFs for the low-rank portion of the model makes subgrid-scale prediction impossible.
@@ 4 Discussion
Neutral temperature fields, especially in the lower thermosphere altitude region, are highly temporally variable because of the influence of atmospheric waves originating from the lower atmosphere. It is helpful to incorporate the dependence of temporal scales on their spatial scales. For this reason, future work could involve extending these ideas into a time-dependent scenario. There are several ways this could be done. Autoregressive structure in the basis coefficients could be endowed, or sets of spatiotemporal functions such as radial 3D needlets could be investigated ([threedneedlets]).
@@ 5 Supplementary Materials
The code and data used in this project are available at https://github.com/mfleduc/FullScaleBGL.
@@ Appendix A Proof of Proposition 1
$L(Q|D,Y)=\log\det\left(\Psi Q^{-1}\Psi^{T}+D\right)+\\ \text{tr}\left(S\left(\Psi Q^{-1}\Psi^{T}+D\right)^{-1}\right)+||\Lambda\circ Q||_{1}$ (23)
@@
$\begin{split}L(Q|D,Y)=\log\det(Q+\Psi^{T}D^{-1}\Psi)+\log\det(D)-\log\det(Q)\\ +\text{tr}\left(SD^{-1}\right)-\text{tr}\left(\Psi^{T}D^{-1}SD^{-1}\Psi\left(Q+\Psi^{T}D^{-1}\Psi\right)^{-1}\right)+||\Lambda\circ Q||_{1}\end{split}$ (26)
@@ Appendix B Derivation of Equation (10)
Starting from the expression of the likelihood given in (9), we note that $\left|\left|\Lambda\circ Q\right|\right|_{1}$ is a convex function of $Q$, and by [convexoptimization] so is $-\log\det(Q)$. By the same logic, $\log\det(Q+\Psi^{T}D^{-1}\Psi)$ is concave in $Q$. Also by [convexoptimization], for positive definite $\Sigma$, $\text{tr}\left(S\Sigma^{-1}\right)$ is convex. So, the problem can be separated into a convex portion given by $f(Q)=-\log\det(Q)+\left|\left|\Lambda\circ Q\right|\right|_{1}$ and a concave portion given by $g(Q)=\log\det(Q+\Psi^{T}D^{-1}\Psi)-\text{tr}\left(\Psi^{T}D^{-1}SD^{-1}\Psi\left(Q+\Psi^{T}D^{-1}\Psi\right)^{-1}\right)$. This gives the likelihood function the form of a difference of two convex functions, more explicitly:
$\begin{split}L(Q|D,Y)=&\underbrace{\log\det(Q+\Psi^{T}D^{-1}\Psi)-\text{tr}\left(\Psi^{T}D^{-1}SD^{-1}\Psi\left(Q+\Psi^{T}D^{-1}\Psi\right)^{-1}\right)}_{concave=g(Q)}\\ &\underbrace{-\log\det(Q)+||\Lambda\circ Q||_{1}}_{convex=f(Q)}\end{split}$ (27)
Now, as done in [bgl], we can apply the Difference of Convex algorithm ([DCA]). To do this we note that since $g(Q)$ is concave, it is bounded above by its linearization $\tilde{g}(Q|Q_{j})$ given by
@@
$\hat{Q}_{j+1}=\underset{Q>0}{\text{argmin}}[\tilde{L}(Q|\hat{Q}_{j},D,Y)]=\underset{Q>0}{\text{argmin}}\left[-\log\det(Q)+\text{tr}\left(\nabla g(\hat{Q}_{j})Q\right)+||\Lambda\circ Q||_{1}\right]$ (30)
@@
[…]

### v2

@@ 4.4.5 Prediction
While the model was fit on data with a $2^{\text{o}}$ resolution in latitude and longitude, the model can be run at higher resolutions, for example we have limited access to $0.25^{\text{o}}$ resolution data. We consider this subgrid-scale data to test the predictive performance of the model by calculating the continuous rank probability score (CRPS) and the root mean square error (RMSE). Reported CRPS and RMSE values are averaged over individual predictions; CRPS has a closed form for Gaussian predictive distribution ([13]),
@@
From these results, it is clear that predictions are much more accurate using the FSBGL with the mean and median CRPS, as well as the RMSE, being less than half that of the BGL and LatticeKrig. In fact, the lowest CRPS obtained by LatticeKrig at any predicted location is 4.64, which is more than the median CRPS for any of the models fitted with the FSBGL. This suggests that the FSBGL is also suitable to be used to fit predictive distributions to data exhibiting multiscale behavior, for example one potential route of future investigation could be whether this technique can be allow for running climate models at lower spatial resolution from which FSBGL can statistically downscale to a fine resolution. This should be especially contrasted with the full-scale approximations of [33], for example, as the reliance on EOFs for the low-rank portion of the model makes subgrid-scale prediction impossible.
@@ 5 Discussion
Neutral temperature fields, especially in the lower thermosphere altitude region, are highly temporally variable because of the influence of atmospheric waves originating from the lower atmosphere. It is helpful to incorporate the dependence of temporal scales on their spatial scales. For this reason, future work could involve extending these ideas into a time-dependent scenario. There are several ways this could be done. Autoregressive structure in the basis coefficients could be endowed, or sets of spatiotemporal functions such as radial 3D needlets could be investigated ([8]). Alternatively, it may be possible to leverage a software package such as EiGLasso ([39]) or TeraLasso ([14]) to determine a Kronecker sum structure using an algorithm similar to the one discussed here.
@@ 6 Supplementary Materials
The code used in this project is available at https://github.com/mfleduc/FullScaleBGL along with some sample datasets.
@@ Appendix A Proof of Proposition 1
[…]
@@
[…]
@@ Appendix B Derivation of Equation (10)
[…]
Now, as done in [20], we can apply the Difference of Convex algorithm ([7]). To do this we note that since $g(Q)$ is concave, it is bounded above by its linearization $\tilde{g}(Q|Q_{j})$ given by
@@
[…]
@@
Appendix C Simulation Studies
C.1 Block-Diagonal model for $Q$
Figure 7: True and estimated $Q$ for the block-diagonal model. The true $Q$ is in the top left, and the remaining figures show the recovered $Q$ with 10, 50, and 100 samples of the field.
Parameter Truth m=10 m=30 m=50 m=100
$Q$ block-diag 0.610 0.444 0.436 0.277
Missed non-zero (%) N/A 51.6 13.8 8.73 0
Missed zero (%) N/A 0.25 4.8 1.54 19.2
$\sigma^{2}$ 1 0.041 0.041 0.059 0.041
$\log_{10}(\tau^{2})$ -2 0.224 0.184 0.232 0.236
$\theta$ 0.3 0.027 0.019 0.043 0.035
$r$ 0.15 0.027 0.043 0.036 0.044
$\nu$ 0.5 0.043 0.073 0.062 0.060
$L(\hat{Q},\hat{p})/L(Q,p)$ 1 1.1 1.1 1.04 1.1
Table 7: Median error in recovery of residual parameters for varying number of independent replicates of the field using a block-diagonal model for $Q$, along with true values. All calculations are the median over 50 trials for a given number of independent replicates of the field. The error in the recovery of $Q$ is given by $\|\hat{Q}-Q\|/\|Q\|$ with $\|\cdot\|$ the Frobenius norm, and the error for the other parameters is the absolute error.
C.2 Hub and Spoke model for $Q$
Figure 8: True and estimated $Q$ for the hub and spoke model. The true $Q$ is in the top left, and the remaining figures show the recovered $Q$ with 10, 50, and 100 samples of the field.
Parameter Truth m=10 m=30 m=50 m=100
$Q$ hub and spoke 0.487 0.336 0.241 0.181
Missed non-zero (%) NA 12.0 1.15 0 0
Missed zero (%) NA 3.34 1.22 8.23 4.1
$\sigma^{2}$ 1 0.059 0.059 0.041 0.059
$\log_{10}(\tau^{2})$ -2 0.342 0.222 0.288 0.270
$\theta$ 0.3 0.035 0.039 0.021 0.027
$r$ 0.15 0.051 0.052 0.045 0.037
$\nu$ 0.5 0.082 0.084 0.071 0.059
$L(\hat{Q},\hat{p})/L(Q,p)$ 1 1.20 1.13 1.12 1.08
Table 8: Median error in recovery of residual parameters for varying number of independent replicates of the field using a hub-and-spoke model for $Q$, along with true values. All calculations are the median over 50 trials for a given number of independent replicates of the field. The error in the recovery of $Q$ is given by $\|\hat{Q}-Q\|/\|Q\|$ with $\|\cdot\|$ the Frobenius norm, and the error for the other parameters is the absolute error.
C.3 Random graph model for $Q$
Figure 9: True and estimated $Q$ for the hub and spoke model. The true $Q$ is in the top left, and the remaining figures show the recovered $Q$ with 10, 50, and 100 samples of the field.
Parameter Truth m=10 m=30 m=50 m=100
$Q$ Random 0.698 0.422 0.315 0.275
Missed non-zero (%) NA 52.8 49.1 32.7 33.5
Missed zero (%) NA 0.877 1.6 6.1 2.57
$\sigma^{2}$ 1 0.241 0.120 0.101 0.601
$\log_{10}(\tau^{2})$ -2 0.284 0.226 0.210 0.200
$\theta$ 0.3 0.213 0.060 0.055 0.108
$r$ 0.15 0.075 0.048 0.052 0.044
$\nu$ 0.5 0.086 0.070 0.054 0.101
$L(\hat{Q},\hat{p})/L(Q,p)$ 1 0.976 0.990 1.04 0.96
Table 9: Median error in recovery of residual parameters for varying number of independent replicates of the field using a random graph model for $Q$, along with true values. All calculations are the median over 50 trials for a given number of independent replicates of the field. The error in the recovery of $Q$ is given by $\|\hat{Q}-Q\|/\|Q\|$ with $\|\cdot\|$ the Frobenius norm, and the error for the other parameters is the absolute error.
C.4 Smaller scale variation
In addition to the residual model explored this far, we also explored the ability to recover a residual model with smaller Matérn scale parameter under the block-diagonal model for $Q$. This greatly limited our ability to recover the taper range, but that parameter is much less important in governing the dynamics of the model in this case, and we are still able to get very near the value of the likelihood at the true parameters. For this reason we do not believe that this is an issue. It is also worth noting that the median error in the estimate for $r$ cannot be any lower due to the resolution of the grid that we are using to do the annealing.
Parameter Truth m=10 m=30 m=50 m=100
$Q$ Block-diagonal 0.632 0.445 0.436 0.278
Missed non-zero (%) NA 52.0 13.8 8.00 0
Missed zero (%) NA 0.223 4.9 1.53 19.4
$\sigma^{2}$ 1 0.081 0.061 0.061 0.061
$\log_{10}(\tau^{2})$ -2 0.278 0.21 0.254 0.290
$\theta$ 0.3 0.675 0.730 0.705 0.699
$r$ 0.05 0.001 0.001 0.001 0.001
$\nu$ 0.5 0.038 0.039 0.039 0.036
$L(\hat{Q},\hat{p})/L(Q,p)$ 1 1.00 0.983 0.992 0.983
Table 10: Median error in recovery of residual parameters with small-scale variation in the residuals using a block-diagonal model for $Q$, along with true values. All calculations are the median over 50 trials for a given number of independent replicates of the field. The error in the recovery of $Q$ is given by $\|\hat{Q}-Q\|/\|Q\|$ with $\|\cdot\|$ the Frobenius norm, and the error for the other parameters is the absolute error.
Appendix D Sample Autocorrelation of the WACCM-X Data
Due to the fact that our samples are output from a time-dependent model, we want to verify that there is limited sample autocorrelation in our data. For this reason, we computed the autocorrelation functions at the spatial locations across samples, a subset of which are plotted here with the corresponding latitude and solar local time. As we can see, the sample autocorrelation is generally small, so we feel comfortable treating these samples as independent.
Figure 10: Sample autocorrelation functions at several latitudes and solar local times (SLT)
Appendix E Behavior of the Nugget Effect
In Section 4.3.1 we hypothesize that the reason that the nugget effect is much smaller for the Tapered Matérn model is that the parameter $\nu$ allows the model to fit the smoothness of the field, while the Gaspar-Cohn and Wendland models do not have the same capability. We provide in this appendix some empirical validation of that assertion.
To investigate the behavior further, we generated 10, 30, 50, and 100 samples of a field on $0,1]^{2}$ using the same cosine basis as the other simulation studies and a block-diagonal precision matrix model. The residual field has the same properties as in C.1, that is, a tapered Matérn covariance with parameters $\tau^{2}=0.01,\sigma^{2}=1,\theta=0.3,r=0.15,\nu=0.5$. However, we are examining whether the Wendland covariance model predicts a larger nugget effect. Our hypothesis is that since the Matérn smoothness is small, leading to a field that is mean-square continuous but not differentiable, and the Wendland covariance model produces smooth fields, we ought to see a much larger nugget effect variance fit in order to compensate. The results of our experiment are shown in Table 11. We see that across sample sizes the estimates of the nugget effect using the Wendland model are greatly inflated, with the fitted nugget standard deviation being 7 times larger under the Wendland model. This supports our hypothesis that the other models may attempt to fit the rougher fields by exaggerating the nugget effect.
Model Truth m=10 m=30 m=50 m=100
Tapered Matérn -2 -2.07 -2.05 -2.01 -2.02
Wendland covariance -2 -0.302 -0.301 -0.306 -0.304
Table 11: Mean estimate of $\log_{10}(\tau^{2})$ on a Matérn residual field with $\tau^{2}=0.01$, for both the Matérn model and a Wendland model. The Wendland model selects a much larger nugget effect, supporting our belief that this might be related to the smoothness of the underlying field.